# Direct numerical simulations of a cylinder cutting a vortex

### v1

@@ V Summary and conclusions
In this manuscript we have conducted direct numerical simulations of normal body-vortex interaction using a thin cylinder. We focused on the transition and the differences in the weak and strong vortex regimes, confirming that it is a smooth transition, primarily driven by the capacity of the cylinder’s boundary layer to detach and wrap around the primary vortex before impact takes place. In the strong vortex regime, this detached, or secondary vorticity propagates along the vortex’s axis, interacting with it and heavily debilitating it, with the result that by the end of the interaction the vortex has almost dissipated. In the weak vortex regime, the cylinder only interacts with the vortex once the body reaches the tube. It then deforms it, propagating waves along the axis, which do not require the modelling of the cylinder’s boundary layer to capture adequately. The behavior of the waves was significantly different to those observed when axial flow is present.
We also investigated the Reynolds number dependence of the flow, showing that while the large contours of behaviour do not change, the vortices left after interaction are much more unstable and show increased turbulence. This also indicates that even in the absence of of background turbulence, the flow left behind the cylinder will become turbulent. These findings both reproduce existing experimental data [22, 17] when the presence of axial flow is not important, and confirm that modelling of the viscous effects are necessary to capture the intricacies of the strong vortex regime. Finally, we measured the force on the cylinder, and showed that while deviations from the baseline values are present, and the force can locally invert direction, the overall force from the interaction is concentrated close to the impact point, and that it averages out on both sides of the interaction.
@@
Our primary interest when simulating normal BVI was on “cutting” vortex lines. However, these simulations showed that the phenomenology is much more complicated, especially for the strong vortex case, and that the interaction deviates from what one could expect from ideal vortex models. While we have examined here normal BVI, with minor modifications the code could be extended to simulating parallel and grazing BVI, and regimes in between. Another possible extension of this work is the question of what happens when a vortex ring- and not a vortex tube is cut. Very few studies of cutting a vortex ring with a wire have been made [4]. Novel structures such as “elephant ears” have been reported, which have not been related to the wider literature on BVI.

### v2

@@ V Summary and conclusions
Results from direct numerical simulations of normal body-vortex interaction using a thin cylinder have been presented. The focus was on the transition and the differences in the weak and strong vortex regimes. Three findings were reported:
  1. 1.
    The simulations confirm that the transition between weak and strong vortex regimes is a smooth transition controlled by $I_{P}$ and primarily driven by the capacity of the cylinder’s boundary layer to detach and wrap around the primary vortex before impact takes place.
  2. 2.
    In the strong vortex regime, this detached, or secondary vorticity propagates along the vortex’s axis, interacting with it and heavily debilitating it, with the result that by the end of the interaction the vortex has almost dissipated.
  3. 3.
    In the weak vortex regime, the cylinder only interacts with the vortex once the body reaches the tube. It then deforms it, propagating waves along the axis. The behavior of the waves was significantly different to those observed when axial flow is present [17].
The Reynolds number dependence of the flow was also investigated by simulating two cases at $Re_{\Gamma}=3000$. The effect of $Re_{\Gamma}$ on the system was reflected in two aspects: at higher $Re_{\Gamma}$ the boundary layers are thinner and have more vorticity, and vortices are more unstable and prone to breaking down. Three main results were found:
  1. 1.
    In the strong vortex regime, qualitative differences in the flow can be seen before the impact of the cylinder. The secondary vorticity, which is now concentrated in smaller and more intense structures, has a stronger interaction with the primary tube leading to more deformations in the neck area. These differences are absent from the weak vortex regime due to the reduced importance of secondary vorticity originating from the boundary layer in the early stages of the interaction.
  2. 2.
    In the weak vortex regime, the stretching and necking of the primary vortex due to the cylinder forcing oppositely signed vorticity close to it is enhanced. However, this does not result in significantly different outcomes.
  3. 3.
    After the interaction, the vortex remnants tend to break down rather than recompose again into a primary tube. The simulations indicate that even in the absence of background perturbations, the flow left behind the cylinder will be turbulent.
The results presented in this manuscript qualitatively reproduce existing experimental data [22, 17] when the presence of axial flow is not important. The simulations also confirm that the adequate modelling of the boundary layer is necessary to capture the intricacies of the strong vortex regime.
The distribution of pressure forces on the cylinder was also reported. The results show local deviations from the baseline value (considered as the one where no vortex is present), which are concentrated around the impact point. The deviations are large, to the point that the force can locally reverse direction, with the cylinder feeling a force in the direction of the motion. However, in general the forces average out on both sides of the interaction, and the integrated force coefficients are very similar to the baseline value.
@@
As the introduction stated, the starting point of the study was on “cutting” vortex lines. However, these simulations showed that the phenomenology is much more complicated, especially for the strong vortex case, and that the interaction deviates from what one could expect from ideal vortex models. While the scope of the manuscript has been restricted to normal BVI, with minor modifications the code could be extended to simulating parallel and grazing BVI, and regimes in between. Another possible extension of this work is the question of what happens when a vortex ring- and not a vortex tube is cut. Very few studies of cutting a vortex ring with a wire have been made [4]. Novel structures such as “elephant ears” have been reported, which have not been related to the wider literature on BVI.
@@ Acknowledgements.
Appendix A Grid dependence studies
To assess the adequateness of the grid, two additional simulations of the case with $I_{P}=0.25$ and $Re_{\Gamma}=3000$ were conducted with coarse and fine grids. These parameters were chosen as this simulation case was the most unfavourable as the cylinder Reynolds number was highest ($Re_{c}=113$), so it provides an upper limit for the problems with the grid. The coarse grid is obtain by halving the resolution in each spatial direction, i.e. $312\times 312\times 936$ points, while the fine grid is obtained by increasing the resolution by a factor 1.5×1.5$\times$1.5 × in every direction, i.e. $936\times 936\times 2808$.
Figure 18 shows the vorticity magnitude for the three grids alongside two plane cuts. The first two rows show the variable at $t\Gamma^{2}/D=220$: this time instant was chosen because it is close to peak dissipation, the tube has been stretched and displaced, and vorticity has intensified which makes it challenging to resolve. Cuts are shown for both $z=0$, i.e. across the wire, $y=0$, i.e. across the vortex core. Clear signatures of underresolution can be seen for the coarse mesh: there are marks of numerical dispersion behind the cylinder in the top-left panel, as well as deformations in the vortex core not seen for the medium and fine grids. While some differences to appear between the medium and fine grids, they are not as significant. This can be quantified by measuring the total circulation in the vortex cores shown in the right column: the coarse grid has a circulation which is $25\%$ lower than the medium grid, while the fine grid has a circulation which is $2\%$ higher than the medium grid. This confirms our assessment that the medium grid can be considered is adequate to within $2\%$, but that further coarsening gives results which are inaccurate due to numerical dispersion.
Finally, the bottom row shows the vorticity magnitude at $t\Gamma^{2}/D=400$ through the $y=0$ plane. This is after the interaction has finished, so the differences between cases should be maximum. Once again, the medium and fine grids show similarities, while the coarse grid shows a picture that is somewhat similar to the other two grids, yet with vorticity concentrated in different points.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 18: Vorticity magnitude for the $I_{P}=0.25$ and $Re_{\Gamma}=3000$ case at $t\Gamma^{2}/D=220$ and $z=0$ (top), at $t\Gamma^{2}/D=220$ and $y=0$ (middle), at $t\Gamma^{2}/D=220$ and $y=0$ (bottom). The left, middle and right columns show the coarse, medium and fine resolutions respectively.
A further check was conducted by assessing the variation in the force on the wire. Figure 19 shows $C_{f}(z)$ at the same time instant for the three meshes. Again, differences can be seen between the three cases, but the difference between the coarse grid and the medium and fine grids is much larger than the difference between medium and fine grids. The integrated force coefficients are $C_{F}=0.24$, $C_{F}=0.31$ and $C_{F}=0.33$ for the coarse, medium and fine grids respectively. This shows that the quantities on the wire are slightly more sensitive than the vorticity statistics. However, the curves all show similar phenomenology, with a strong suction peak, and a strong increase close to the vortex core. Hence, we do not expect the results from the medium mesh to miss any important physics, while the coarse mesh shows $C_{f}$ dropping below the baseline value, questioning its use.
To summarize, the medium mesh shows small discrepancies when compared to a finer mesh, especially when measuring the force on the wire. However, the overall phenomenology is well captured, unlike what is seen for the coarser mesh, where the flow topology is changed. We also reiterate that this is the most unfavourable instance, and yet the medium grid is able to produce reasonable results with a much lower computational cost.
Figure 19: Force coefficient at $t\Gamma^{2}/D=220$ for the $I_{P}=0.25$ and $Re_{\Gamma}=3000$ case for the three grids. A horizontal line shows the baseline value when no vortex is present. Symbols: blue, coarse grid; orange, medium grid; green, fine grid.